OpenStudy (anonymous):

2^5/2-2^3/2 doesnt it equal 2^2

OpenStudy (anonymous):

or just 2?

OpenStudy (mertsj):

\[\frac{2^5}{2^1}- \frac{2^3}{2^1}=2^{5-1}-2^{3-1}=2^4-2^2=16-4=12\]

OpenStudy (mertsj):

No!!! @mathstudent55 Only like terms can be added. Like terms have the same exponents.

OpenStudy (anonymous):

\[2^{5/2}-2^{3/2}\] answers are 2^1/2,2,2^3/2,2^5/3,2^2 it says the correct answers is 2^3/2

OpenStudy (anonymous):

i dont get how this is possible, here is the test it is the first in college level algebra http://media.collegeboard.com/digitalServices/pdf/accuplacer/accuplacer-sample-questions-for-students.pdf

OpenStudy (anonymous):

*level mathmatics

OpenStudy (mathstudent55):

@Mertsj Wow. Midway through the problem I forgot it was a subtraction instead of a division. I still think the question is using fractional exponents, though.

OpenStudy (anonymous):

it is fractional exponents

OpenStudy (mertsj):

\[2^{\frac{5}{2}}=\sqrt{2^5}=\sqrt{32}=4\sqrt{2}\]

OpenStudy (mertsj):

\[2^{\frac{3}{2}}=\sqrt{2^3}=\sqrt{8}=2\sqrt{2}\]

OpenStudy (mertsj):

\[4\sqrt{2}-2\sqrt{2}=2\sqrt{2}\]

OpenStudy (anonymous):

i really dont understand how they say the correct answer is \[2^{3/2}\]

OpenStudy (mertsj):

\[2\sqrt{2}=2^1(2)^{\frac{1}{2}}=2^{1+\frac{1}{2}}=2^{\frac{3}{2}}\]

OpenStudy (mertsj):

Here is another way to think of it:

OpenStudy (anonymous):

where did you get? [2^{1}\]? sorr this is so confusing

OpenStudy (mertsj):

\[2^{\frac{5}{2}}-2^{\frac{3}{2}}=2^{\frac{3}{2}}(2^{\frac{2}{2}}-1)=2^{\frac{3}{2}}(2-1)=2^{\frac{3}{2}}(1)=2^{\frac{3}{2}}\]

OpenStudy (mertsj):

2 is the same as 2^1

OpenStudy (mathstudent55):

\( \large{ 2^{\frac{5}{2}}-2^{\frac{3}{2} } }\) \( = 2 \times 2^{\frac{3}{2}} - 2^{\frac{3}{2}} \) \(= 2^{\frac{3}{2}} (2 - 1) \) \(= 2^{\frac{3}{2}} \)

OpenStudy (mertsj):